solve
11w - 11 = -77 .

Answer by maswartz(2)

11w-11=-77 Isolate the w to one side by moving the negative 11 over to the other side. It becomes postive when it moves. 11w=-77+ 11 or 11w=-66 11*6 is 66 so your answer is w=-6.